### v1

@@ 3.4 Collaborative Video Generating Branch
Feature Integration Layer. Inspired by the prior work [120, 119], we adopt Adaptive Instance Normalization (AdaIN) for feature fusion. We firstly align the dimensions of the dynamic motion feature $f_{d}$ and static context feature $f_{s}$ by MLPs. The fusion of $f_{d}$ and $f_{s}$ is accomplished through a denormalization process. The detailed structure of the feature integration layer is shown in Fig. 2.
@@
$\begin{split}\min_{G^{r}_{v}}&\max_{D^{r}_{v}}\mathcal{L}_{r}=\\ &E_{V_{r}}[\log D^{r}_{v}(V_{r})]+E_{z}[\log(1-D^{r}_{V}(G^{r}_{V}(z^{r}_{v}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{r}(h,w)-g_{t}^{r}(h,w)||_{2}^{2}\end{split}$ (6)
$\begin{split}\min_{G^{p}_{v}}&\max_{D^{p}_{v}}\mathcal{L}_{p}=\\ &E_{V_{p}}[\log D^{p}_{v}(V_{p})]+E_{z}[\log(1-D^{p}_{V}(G^{p}_{V}(z_{v}^{p}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{p}(h,w)-g_{t}^{p}(h,w)||_{2}^{2}\end{split}$ (7)
@@
where $V_{r}$ and $V_{p}$ denote the generation target of each generator videos and $T$, $H$, $W$ represent the temporal and spatial dimensions of video clip, respectively. $v_{t}^{r}(h,w)$ and $v_{t}^{p}(h,w)$ indicates the pixel value at the $(h,w)$ coordinates on the $t$-th frame in the ground truth video. The generated videos are denoted as $g_{t}^{r}(h,w)$ and $g_{t}^{p}(h,w)$. Similar to pose prediction branch, we use the OSGANs parameter updating strategy [50] to simplify the training process. Two parallel training objectives enhance the understanding of the temporal relationship of the video, leading to a more comprehensive video representation.
3.5 Joint Optimization
@@
where the $\sigma_{p},\sigma_{c}$ are scalar hyper-parameters. All the networks in three branches are jointly optimized with $\mathcal{L}$ in an end-to-end fashion. Finally, for fine-tuning on downstream tasks, we utilize the pretrained parameters from the V-Network as the initialization.
4 Experiments
In this section, we conduct experiments on three popular action recognition datasets, i.e., UCF101 [121], HMDB51 [122], and Kinetics400[148] to verify the effectiveness of our CSVR. Besides, we also conduct ablation studies to investigate the significance of each branch in our method.
4.1 Dataset
The experiments of CSVR are mainly conducted on three datasets. UCF101 contains 13,320 videos covering 101 action classes, while HMDB51 consists of 6,766 videos in 51 classes. Following previous methods [94, 160], we use split 1 of UCF101 and HMDB51 in our experiments. Kinetics400 is a large-scale dataset containing 246K videos in its training set which includes 400 classes of human actions. During the experiments, we pretrain our CSVR on UCF101 training set and Kinetics400 training set, respectively. Then we finetune the pretrained V-Network to two downstream tasks, i.e., action recognition and video retrieval, on UCF101 and HMDB51, respectively.
4.2 Implementation Details
Backbone For context matching branch, we explore four 3D-CNN backbones as V-Network, i.e., C3D [149], R(2+1)D[150], R3D [151, 80], and S3D [152] to extract video representations. Moreover, we adopt R2D-10 as I-Network to extract features from I-frame. For pose prediction branch, we utilize 1D-CNN-based CGAN for pose prediction. In addition, R3D-based TGAN v2 is leveraged in video generating branch. It should be noted that during inference, the major computation cost is from video representation extraction V-Network which is mainly determined by various 3D-CNN backbones.
@@ 4.2 Implementation Details
As for the generation of $f_{i}$, the feature of the last layer in pose generator is utilized as the the dynamic motion feature $f_{d}$. After investigating the selection of hyper parameters, which is introduced in Section 4.3, we set $\sigma_{p}=\sigma_{c}=0.75$ in our final experiment.
Finetuning After pretraining our CSVR on video dataset, the integration feature $f_{i}$ from the V-Network is utilized to finetune on downstream tasks. We conduct experiments on two widely-used human action datasets, i.e., UCF101 and HMDB51, and report Top-1 accuracy of action recognition task and Recall at k (R@k) of video retrieval task on each dataset.
For action recognition task, we initialize the backbone with the pretrained parameters except for the last fully connected layer. We finetune the model with a batch size $B$ of 32 and use the cosine annealing scheduler to decay the learning rate during training. For UCF101, we set the learning rate to $lr=0.0001\times B$ and the weight decay of the SGD optimizer to 0.003. For HMDB51, we set the learning rate to $lr=0.0002\times B$ and the weight decay to 0.002. We incorporate a dropout ratio of 0.3 in the second-to-last layer of the network during finetuning. Data augmentation methods, e.g., random crop, random flip, Gaussian blur, and color jitter are applied to increase the diversity of the training data.
@@
4.3 Ablation Study
To further investigate the effectiveness of each branch of CSVR, we conduct experiments by firstly pretraining on UCF101 training set and then finetuning on the same set for action recognition. We choose R(2+1)D as the backbone of V-network for all ablation study experiments.
Table 1: Action recognition performance comparison of different branches. 4-th row represents the integrated feature $f_{i}$ without optimization from video generating branch, while in the last row, $f_{i}$ is learned by the full CSVR pretraining framework.
$f_{d}$ $f_{s}$ $f_{i}$ $V$ UCF101 HMDB51
@@ 4.3 Ablation Study
90.3 56.5
Table 2: Correlation between pretext and downstream tasks. The linear relationship between the results indicates the effectiveness of our pretraining. MSE score is increased by a factor of 100 to emphasize the change. All the experiments are conducted on UCF101 dataset.
@@
Effectiveness of Different Branches. Since our method includes three different branches, we evaluate the contribution of each branch to the performance improvement. Table 1 shows the final action recognition performance on UCF101 and HMDB51 with features from various branches as the initialization for finetuning on downstream action recognition. For inference, we transfer different latent features to V-Network and finetune them on downstream tasks. As seen, both the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ significantly improve the action recognition accuracy compared to the baseline of training from scratch, which demonstrate the superiority of designing context matching and pose prediction branches. Interestingly, the dynamic motion feature $f_{d}$ achieves better results than static context feature $f_{s}$ on HMDB51, while $f_{d}$ underperform $f_{s}$ on UCF101. This may be because most classes in UCF101 can be classified by context information, while classes in HMDB51 are mainly distinguished by the motion information. Besides, we achieve the integrated feature $f_{i}$ by fusing the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ with AdaIN. By employing $f_{i}$ for the downstream task finetuning, the improvement on accuracy further illustrates the complementarity of static context and dynamic motion features for action recognition. Finally, we introduce video generating branch to further improve the integrated feature $f_{i}$ by jointly optimizing three branches. The accuracy on UCF101 and HMDB51 is significantly improved to 90.3% and 56.5%, respectively, which shows the effectiveness of collaboratively learning static context and dynamic motion features for action recognition.
Correlation between pretext and downstream tasks. We conduct experiments to explore the correlation between the pretext tasks and the downstream task by measuring their performance during different pretraining epochs. The performance measurements include the top-1 accuracy of context matching, mean squared error (MSE) of pose prediction, inception score (IS) of generated videos and top-1 accuracy of downstream action recognition. The results in Table 2 show a consistent improvement in action recognition accuracy as the performance of the pretext tasks increases, which indicates the strong correlations between various pretext and downstream tasks.
Table 3: Comparison of $f_{d}$ from different layers of $G_{p}$ encoder.
Table 4: Comparison of $f_{s}$ with different loss functions.
@@
Table 4: Comparison of $f_{s}$ with different loss functions.
Selection of Dynamic Motion Feature. In order to locate the hidden layer which enriches the most dynamic information, we pretrain the pose generator $G_{p}$ on poses from UCF101 dataset and use features from different layers of encoder in $G_{p}$ as dynamic motion features $f_{d}$. We report the results of different $f_{d}$s finetuned on action recognition task. As shown in Table 4, $f_{d}$ from the last layer achieves the best performance.
@@
$\begin{split}\mathcal{L}_{context}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})}{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x_{i})}{\tau})}\end{split}$ (10)
@@
where $x_{i}$ denotes the global video representation, and $z_{i}^{*}$ and $z_{k}^{*}$ denote the positive and negative I-frame features, other symbols are the same as Equation 2. However, there may exist multiple I-frames in a single video clip, indicating more positive samples. Hence, we apply more advanced Multi-Instance-InfoNCE (MI-InfoNCE) [161] as the contrast loss in static context branch. Comparing to the traditional InfoNCE loss utilized in [160], MI-InfoNCE loss enables more adequate comparison between positive and negative samples, resulting in better performance on downstream tasks. To further verify the effectiveness of contrastive loss, we also conduct ablation experiments with L2 loss. When calculating InfoNCE and L2 loss, we randomly select an I-frame from the video clip as learning objective to ensure that the loss function produces correct outputs. As shown in Table 4, MI-InfoNCE achieve optimal results among three losses. Interestingly, we observe that both InfoNCE and MI-InfoNCE achieve identical scores when testing on HMDB51, while MI-InfoNCE outperforms the others on UCF101. This is due to action recognition on UCF101 relies more on static context information.
@@
Selection of Video Generation Target Since our video generating branch is trained with two different objective functions, we present the performance of each separately on the downstream tasks. Therefore, we train models in three different setups, i.e., jointly train CSVR with both $\mathcal{L}_{r}$ and $\mathcal{L}_{p}$, and train CSVR with $\mathcal{L}_{r}$ or $\mathcal{L}_{p}$ independently. As shown in Table 6, the results obtained from the full branch objective surpass the result of reconstruction objective in both cases. Moreover, when compared to the result of prediction objective, the full branch objective outperforms better on the UCF101 dataset and performs equally well on HMDB51.
Hyperparameter investigation of $\sigma_{p},\sigma_{c}$. CSVR contains three branches which are jointly optimized by loss function $L$. $\sigma_{p},\sigma_{c}$ in $L$ are the hyperparameters that coordinate the weights of pose prediction branch and context matching branch, respectively. We perform a grid search on all combinations of two parameters within a range of 0.5 - 1.5 and conduct the pretraining experiments with each combination to find the optimal parameters. Figure 6 shows the results on UCF101 when transferring the pretrained features to the downstream action recognition task. As seen, the best results are achieved when $\sigma_{p}=\sigma_{c}=0.75$, which are chosen for all other experiments.
@@
Table 5: Performance comparison of using different objective functions in video generating branch to train CSVR.
Objective UCF101 HMDB51
$\mathcal{L}_{r}$ 89.2 54.6
$\mathcal{L}_{p}$ 89.9 56.6
$\mathcal{L}_{r}$ $\mathcal{L}_{p}$ 90.3 56.5
@@
Table 6: Action recognition results on UCF101 with different combinations of $\sigma_{p},\sigma_{c}$.
@@
4.4 Comparison with State-of-the-art Methods
To verify the effectivenss of our CSVR framework, we conduct experiments on two downstream tasks, i.e., action recognition and video retrieval, with four 3D-CNN backbones, i.e., R(2+1)D, R3D, C3D and S3D. The whole model is pretrained on two datasets, i.e., UCF101 and large-scale Kinetics400, respectively.
Table 7: Action Recognition. Comparison of CSVR with state-of-the-art self-supervised methods on the action recognition task. *STCNet is a specially-designed R3D network used in DynamoNet[162].
Method Backbone Pretrain Dataset Epochs Resolution Clip Length UCF101 HMDB51

### v2

@@ III-D Collaborative Video Generating Branch
Feature Integration Layer. Inspired by the prior work [119, 118], we adopt Adaptive Instance Normalization (AdaIN) for feature fusion. We firstly align the dimensions of the dynamic motion feature $f_{d}$ from the pose prediction branch and static context feature $f_{s}$ from V-network by adding convolutional layers and MLPs after the encoder of $G_{p}$. Then, the fusion of $f_{d}$ and $f_{s}$ is accomplished through a denormalization process. The detailed structure of the feature integration layer is shown in Fig. 2.
@@
$\begin{split}\min_{G^{r}_{v}}&\max_{D^{r}_{v}}\mathcal{L}_{r}=\\ &\mathbb{E}_{v\sim V_{r}}[\log D^{r}_{v}(V_{r})]+\mathbb{E}_{z\sim p(z)}[\log(1-D^{r}_{V}(G^{r}_{V}(z^{r}_{v}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{r}(h,w)-g_{t}^{r}(h,w)||_{2}^{2}\end{split}$ (6)
$\begin{split}\min_{G^{p}_{v}}&\max_{D^{p}_{v}}\mathcal{L}_{p}=\\ &\mathbb{E}_{v\sim V_{p}}[\log D^{p}_{v}(V_{p})]+\mathbb{E}_{z\sim p(z)}[\log(1-D^{p}_{V}(G^{p}_{V}(z_{v}^{p}|f_{i}))]+\\ &\frac{1}{THW}\sum^{THW}||v_{t}^{p}(h,w)-g_{t}^{p}(h,w)||_{2}^{2}\end{split}$ (7)
@@
where $V_{r}$ and $V_{p}$ denote the generation target of each generator videos and $T$, $H$, $W$ represent the temporal and spatial dimensions of video clip, respectively. $v_{t}^{r}(h,w)$ and $v_{t}^{p}(h,w)$ indicate the pixel value at the $(h,w)$ coordinates on the $t$-th frame in the ground truth video. The generated videos are denoted as $g_{t}^{r}(h,w)$ and $g_{t}^{p}(h,w)$. Similar to pose prediction branch, we use the OSGANs parameter updating strategy [50] to simplify the training process. Two parallel training objectives enhance the understanding of the temporal relationship of the video, leading to a more comprehensive video representation.
III-E Joint Optimization
@@
where the $\sigma_{p},\sigma_{c}$ are scalar hyper-parameters. Since our CSVR contains three branches, we firstly pretrain each branch separately and then jointly optimize them with $\mathcal{L}$ in an end-to-end fashion for stable training. Finally, for fine-tuning on downstream tasks, we utilize the pretrained integrated feature $f_{i}$ as the initialization.
IV Experiments
In this section, we conduct experiments on four popular action recognition datasets, i.e.UCF101 [120], HMDB51 [121], Something-Something v2(SSv2) [169] and Kinetics-400 [148] to verify the effectiveness of our CSVR. Besides, we also conduct ablation studies to investigate the significance of each branch in our method.
IV-A Dataset
The experiments of CSVR are mainly conducted on four datasets. UCF101 contains 13,320 videos covering 101 action classes, while HMDB51 consists of 6,766 videos in 51 classes. Following previous methods [94, 160], we use split 1 of UCF101 and HMDB51 in our experiments. SSv2 is another extensive video dataset, emphasizes temporal information with its 169K training and 20K validation videos across 174 action classes. Kinetics-400 is a large-scale dataset containing 246K videos in its training set which includes 400 classes of human actions. During the experiments, we pretrain our CSVR on UCF101 training set and Kinetics-400 training set, respectively. Then we finetune the network initialized with the pretrained $f_{i}$ to two downstream tasks, i.e., action recognition and video retrieval, on UCF101 and HMDB51, respectively.
IV-B Implementation Details
Backbone For context matching branch, we explore four 3D-CNN backbones as V-Network, i.e., C3D [149], R(2+1)D[150], R3D [151, 80], and S3D [152] to extract video representations. Moreover, we adopt R2D-10 as I-Network to extract features from I-frame. For pose prediction branch, we utilize 1D-CNN-based CGAN for pose prediction. In addition, R3D-based TGAN v2 is leveraged in video generating branch. It should be noted that during inference, the major computation cost is from video representation extraction network which is mainly determined by various 3D-CNN backbones.
@@ IV-B Implementation Details
As for the generation of $f_{i}$, the feature of the last layer in pose generator is utilized as the the dynamic motion feature $f_{d}$. After investigating the selection of hyper parameters, which is introduced in Section IV-C, we set $\sigma_{p}=\sigma_{c}=0.75$ in our final experiment.
Finetuning After pretraining our CSVR on video dataset, the integrated feature $f_{i}$ is utilized to finetune on downstream tasks. We conduct experiments on two widely-used human action datasets, i.e., UCF101 and HMDB51, and report Top-1 accuracy of action recognition task and Recall at k (R@k) of video retrieval task on each dataset.
For action recognition task, we initialize the backbone with the pretrained parameters except for the last fully connected layer. We finetune the model with a batch size $B$ of 32 and use the cosine annealing scheduler to decay the learning rate during training. For UCF101, we set the learning rate to $lr=0.0001\times B$ and the weight decay of the SGD optimizer to 0.003. For HMDB51, we set the learning rate to $lr=0.0002\times B$ and the weight decay to 0.002. Pretrain models are finetuned on both datasets for 100 epochs. We incorporate a dropout ratio of 0.3 in the second-to-last layer of the network during finetuning. Data augmentation methods, e.g., random crop, random flip, Gaussian blur, and color jitter are applied to increase the diversity of the training data.
@@
IV-C Ablation Study
To further investigate the effectiveness of each branch of CSVR, we conduct experiments by firstly pretraining on UCF101 training set and then finetuning on the same set for action recognition. We choose R(2+1)D as the backbone of V-network and finetuning network for all ablation studies.
TABLE I: Action recognition performance comparison of different branches. In table below, $f_{i}$(AdaIN) and $f_{i}$(Concate) represents the integrated feature $f_{i}$ generated by AdaIN or directly concatenating, both of which are generated without optimization from video generating branch. $f_{i}$(full) indicates $f_{i}$ learned by the full CSVR pretraining framework.
$f_{d}$ $f_{s}$ $f_{i}$(AdaIN) $f_{i}$(Concat) $f_{i}$(full) UCF101 HMDB51
@@ IV-C Ablation Study
75.2 41.7
90.4 56.6
Effectiveness of Static and Dynamic Features. Since both static information and dynamic motion are crucial to understanding actions, we evaluate the contribution of features from the pose prediction branch and the context matching branch to the performance improvement in this experiment. Table I shows the final action recognition performance on UCF101 and HMDB51 with features from various branches as the initialization for finetuning on downstream action recognition. As seen, both the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ significantly improve the action recognition accuracy compared to the baseline of training from scratch, demonstrating the superiority of designing context matching and pose prediction branches. Interestingly, the dynamic motion feature $f_{d}$ achieves better results than static context feature $f_{s}$ on HMDB51, while $f_{d}$ underperform $f_{s}$ on UCF101. This may be because most classes in UCF101 can be classified by context information, while classes in HMDB51 are mainly distinguished by the motion information.
Effectiveness of Video Generating Branch. The collaborative video generating branch contains two components: the integration layer and the video generating networks. Following previous practice, we evaluate the effectiveness of these components by finetuning the feature on action recognition tasks on UCF101 and HMDB51. By default, the integrated feature $f_{i}$ is generated by fusing the static context feature $f_{s}$ and dynamic motion feature $f_{d}$ with AdaIN. As seen in Table. I, when employing $f_{i}$ for the downstream task finetuning, the improvement on accuracy further illustrates the complementarity of static context and dynamic motion features for action recognition. To further emphasize the effectiveness of AdaIN, we directly concatenate $f_{d}$ and $f_{s}$, followed by convolution to generate the concat feature $f_{i}$(Concat). It can be observed that the concat feature offers slight improvements over those produced by previous branches in downstream tasks and performs much worse than the feature generated by AdaIN. Finally, we use the full CSVR framework to further improve the integrated feature $f_{i}$ by jointly optimizing three branches. The accuracy on UCF101 and HMDB51 is significantly improved to 90.3% and 56.5%, respectively, which shows the effectiveness of collaboratively learning static context and dynamic motion features for action recognition.
TABLE II: Correlation between pretext and downstream tasks. The linear relationship between the results indicates the effectiveness of our pretraining. MSE score is increased by a factor of 100 to emphasize the change. All the experiments are conducted on UCF101 dataset.
@@
Correlation between pretext and downstream tasks. We conduct experiments to explore the correlation between the pretext tasks and the downstream task by measuring their performance during different pretraining epochs. The performance measurements include the top-1 accuracy of context matching, mean squared error (MSE) of pose prediction, inception score (IS) of generated videos and top-1 accuracy of downstream action recognition. The results in Table II show a consistent improvement in action recognition accuracy as the performance of the pretext tasks increases, which indicates the strong correlations between various pretext and downstream tasks.
TABLE III: Comparison of feature $f_{d}$ from different layers of $G_{p}$ encoder.
@@
TABLE IV: Comparison of feature $f_{s}$ with different loss functions.
@@
Selection of Dynamic Motion Feature. In order to locate the hidden layer which enriches the most dynamic information, we pretrain the pose generator $G_{p}$ on poses from UCF101 dataset and use features from different layers of encoder in $G_{p}$ as dynamic motion features $f_{d}$. We report the results of different $f_{d}$ finetuned on action recognition task. As shown in Table IV, $f_{d}$ from the last layer achieves the best performance.
@@
$\begin{split}&\mathcal{L}_{context}=\\ &-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})}{\exp(\frac{\cos(z^{*}_{i},x_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x_{i})}{\tau})}\end{split}$ (10)
@@
where $x_{i}$ denotes the global video representation, and $z_{i}^{*}$ and $z_{k}^{*}$ denote the positive and negative I-frame features, other symbols are the same as Equation 2. However, there may exist multiple I-frames in a single video clip, indicating more positive samples. Hence, we apply more advanced Multi-Instance-InfoNCE (MI-InfoNCE) [161] as the contrast loss in static context branch. Comparing to the traditional InfoNCE loss utilized in [160], MI-InfoNCE loss enables more adequate comparison between positive and negative samples, resulting in better performance on downstream tasks. To further verify the effectiveness of contrastive loss, we also conduct ablation experiments with L2 loss. When calculating InfoNCE and L2 loss, we randomly select an I-frame from the video clip as learning objective to ensure that the loss function produces correct outputs. As shown in Table IV, MI-InfoNCE achieve optimal results among three losses. Interestingly, we observe that both InfoNCE and MI-InfoNCE achieve identical scores when testing on HMDB51, while MI-InfoNCE outperforms the others on UCF101. This is due to action recognition on UCF101 relies more on static context information.
@@
Selection of Video Generation Target Since our video generating branch is trained with two different objective functions, we present the performance of each separately on the downstream tasks. Therefore, we train models in three different setups, i.e., jointly train CSVR with both $\mathcal{L}_{r}$ and $\mathcal{L}_{p}$, and train CSVR with $\mathcal{L}_{r}$ or $\mathcal{L}_{p}$ independently. As shown in Table V, the results obtained from the full branch objective surpass the result of reconstruction objective in both cases. Moreover, when compared to the result of prediction objective, the full branch objective outperforms better on the UCF101 dataset and performs equally well on HMDB51. When conducting experiments on the SSv2 dataset, which is more sensitive to temporal information, the performance using the full branch objective significantly surpassed that of using $\mathcal{L}_{r}$ or $\mathcal{L}_{p}$ independently.
@@
TABLE V: Performance comparison of using different objective functions in video generating branch to train CSVR.
Objective UCF101 HMDB51 SSv2
$\mathcal{L}_{r}$ 89.2 54.6 56.1
$\mathcal{L}_{p}$ 89.9 56.6 55.3
$\mathcal{L}_{r}$ $\mathcal{L}_{p}$ 90.4 56.6 59.2
@@
Figure 3: Action recognition results on UCF101 with different combinations of $\sigma_{p},\sigma_{c}$.
Hyperparameter Investigation of $\sigma_{p},\sigma_{c}$. CSVR contains three branches which are jointly optimized by loss function $L$. $\sigma_{p},\sigma_{c}$ in $L$ are the hyperparameters that coordinate the weights of pose prediction branch and context matching branch, respectively. We perform a grid search on all combinations of two parameters within a range of 0.5 - 1.25 and conduct the pretraining experiments with each combination to find the optimal parameters. Figure 3 shows the results on UCF101 when transferring the pretrained features to the downstream action recognition task. As seen, the best results are achieved when $\sigma_{p}=\sigma_{c}=0.75$, which are chosen for all other experiments.
@@
IV-D Comparison with State-of-the-art Methods
To verify the effectivenss of our CSVR framework, we conduct experiments on two downstream tasks, i.e., action recognition and video retrieval, with four 3D-CNN backbones, i.e., R(2+1)D, R3D, C3D and S3D. The whole model is pretrained on two datasets, i.e., UCF101 and large-scale Kinetics-400, respectively.
TABLE VI: Action Recognition on UCF101 and HMDB51. Comparison of CSVR with state-of-the-art self-supervised methods on the action recognition task. *STCNet is a specially-designed R3D network used in DynamoNet[162].
{tabu}
lc—cccc—cc Method Backbone Pretrain Dataset Epochs Resolution Clip Length UCF101 HMDB51